Suppose there are lemon (used) cars worth every price from ZMW1,000 to ZMW3000 to new owners, with each price equally likely. Further you expect that a car worth X to a new owner is worth 0.75X to its current owner. Suppose that there is a finite supply of cars at each price, while there is a much larger number of potential buyers (so the equilibrium price is the valuation of the potential buyer). Lastly assume that the quality of the car is unknown to potential buyers, due to information asymmetry.

(a) Calculate the equilibrium price and average quality of cars sold.


(b) Which cars are sold and which ones are not?


(c) Is the equilibrium efficient?


1
Expert's answer
2021-09-02T10:47:02-0400

a)

Each price have a probability of "\frac{1}{2}" due to information asymmetry.

Average price =probabilities of car

"=(\frac{1}{2}\times 1000)+(\frac{1}{2}\times 3000)=2000"

The equilibrium price will be ZMW 2000


b)

Average price is ZMW 2000

The worth of a car is 0.75 its value.

"0.75\times3000=2250"

The buyers are willing to pay the average price thus the car price of ZMW 2250 wont be sold.


c)

The equilibrium is not efficient. A market with information asymmetry generate surplus and loss.